In 2019, Lisa and Fred, a married couple, had taxable income of $300,000. If they were to file separate tax returns, Lisa would have reported taxable income of $125,000 and Fred would have reported taxable income of $175,000. Use Tax Rate Schedule for reference.

What is the couple’s marriage penalty or benefit? (Do not round intermediate calculations.)

2019 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$           0 $ 19,400 10% of taxable income
$ 19,400 $ 78,950 $1,940 plus 12% of the excess over $19,400
$ 78,950 $168,400 $9,086 plus 22% of the excess over $78,950
$168,400 $321,450 $28,765 plus 24% of the excess over $168,400
$321,450 $408,200 $65,497 plus 32% of the excess over $321,450
$408,200 $612,350 $93,257 plus 35% of the excess over $408,200
$612,350 $164,709.50 plus 37% of the excess over $612,350

Schedule Z-Head of Household

If taxable income is over: But not over: The tax is:
$           0 $ 13,850 10% of taxable income
$ 13,850 $ 52,850 $1,385 plus 12% of the excess over $13,850
$ 52,850 $ 84,200 $6,065 plus 22% of the excess over $52,850
$ 84,200 $160,700 $12,962 plus 24% of the excess over $84,200
$160,700 $204,100 $31,322 plus 32% of the excess over $160,700
$204,100 $510,300 $45,210 plus 35% of the excess over $204,100
$510,300 $152,380 plus 37% of the excess over $510,300

Schedule Y-2-Married Filing Separately

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $306,175 $46,628.50 plus 35% of the excess over $204,100
$306,175 $82,354.75 plus 37% of the excess over $306,175
Answer #1
Tax payable, if filed jointly 60349.00 =28765+(300000-168400)*24%
Tax payable, if filed single:
Lisa 24174.5 =14382.5+(125000-84200)*24%
Fred 37316.5 =32748.5+(175000-160725)*32%
Tax payable, if filed single 61491.00
Couple’s marriage benefit 1142.00 =61491-60349